electric potential is given by :V=6x-8xy^2-8y+6yz-4z^2 then the electric force acting on 2 coulomb point charge placed on origin will be:                (1) 2N    (2) 6N     (3)8N     (4) 20N     

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
As we know
 
Ex = -dV /dx = -6 + 8y²
Similarly
Ey = -dV/dy = 16xy +8 - 6z
 
Ez = -6y + 8z
 
If chrge is at (0,0,0)
Then
 
Ex = -6
Ey = 8
Ez = 0
 
Hence
Enet = sqrt(6²+8²) = 10
 
Now
 
F = q *Enet
 
Then
F = 2 * 10 = 20 N